Circular No. 7360

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000F IN IC 302 P. Garnavich, University of Notre Dame; and S. Jha, P. Challis, and

R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of 2000F (cf. IAUC 7353) was obtained by E. Falco on Feb. 2.1 UT with the F. L. Whipple Observatory 1.5-m telescope. The spectrum was taken under poor conditions and is of low signal-to-noise but suggests that the supernova is of type Ic. The continuum is red; a broad absorption is seen at a rest wavelength of 570 nm and a weak emission band is visible at 460 nm. Narrow Hα emission from the host galaxy provides a recession velocity of 5850 km/s.

Preliminary CCD photometry of SN 2000F by K. Krisciunas, University of Washington, and F. Deglman, Apache Point Observatory (APO), using the APO 3.5-m telescope: Feb. 4.09 UT, V = 18.69, B − V = +1.88, V − R = +0.78, V − I = +1.33 (uncertainties roughly ± 0.03); Feb. 7.09, 18.87 ± 0.02, +1.98 ± 0.06, +0.84 ± 0.02, +1.45 ± 0.02.

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD Garnavich, Kirshner, and Challis, on behalf of the Supernova Inten-

sive Study team, also report that Hubble Space Telescope (HST) Wide Field/Planetary Camera 2 (WFPC2) images of SN 1987A were obtained on Feb. 2 UT, which show that additional bright spots have developed around the inner ring since the previous HST imaging in 1999 April. The ‘structure’ on the southeast quadrant of the ring reported by Bouchet et al. (IAUC 7354) is resolved into four individual spots that appear unresolved at the WFPC2 resolution. An additional spot is developing in the southwest quadrant, and there appear to be three faint spots forming on the northwest side, but these are at the limit of detectability. The original region of activity (‘spot 1’; see below) continues to brighten, but at a slower rate than previously seen (IAUC 7102). The distance and p.a. from the supernova position and Hα fluxes (in units of 10^{−15} erg cm^{−2} s^{−1}) for each spot are as follows: spot 1, 0′′.58 in p.a. 29^{o}, 13.8; spot 2, 0′′.73 in p.a. 91^{o}, 1.3; spot 3, 0′′.70 in p.a. 106^{o}, 1.6; spot 4, 0′′.63 in p.a. 122^{o}, 2.0; spot 5, 0′′.58 in p.a. 139^{o}, 0.4; spot 6, 0′′.66 in p.a. 229^{o}, 0.9. The current Hα flux of spot 4 is similar to that of spot 1 in early 1997. Spot 6 is within 0′′.15 of a 20th-magnitude star projected on the ring (Plait et al. 1995, Ap.J. 439, 730). The sudden increase in activity suggests that the full-scale collision between the supernova ejecta and the circumstellar ring gas has finally begun (e.g., McCray 1993, Ann. Rev. A.Ap. 31, 175).

2000 February 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green